OpenStudy (anonymous):

The Value of 2[ ( cos 50 + i sin50) * (cos40 * i cos40)} is?

OpenStudy (anonymous):

@mukushla @ghazi @UnkleRhaukus @lgbasallote @eliassaab @lalaly

OpenStudy (ghazi):

use De movire's theorem

OpenStudy (anonymous):

Euclid

OpenStudy (ghazi):

moivre*

OpenStudy (anonymous):

e^itheata

OpenStudy (anonymous):

\[e^i \theta\]

OpenStudy (anonymous):

First CORRECTION is probably \[2[(\cos 50 + i \sin 50) *(\cos40 + i \sin40)]\]

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

Then use indeed Euler-DeMoivre formula\[\cos \theta + i \sin \theta = e^{i \theta}\]

OpenStudy (anonymous):

thxxx

OpenStudy (anonymous):

2i

OpenStudy (ghazi):

\[e^{i50}*e^{i40}\]

OpenStudy (anonymous):

NO you made a mistake

OpenStudy (ghazi):

calculate (cos90+isin90)

OpenStudy (anonymous):

i got itthxx

OpenStudy (anonymous):

2i